OpenStudy (anonymous):

Polynomial Functions Evaluate f(2t-1)= (2t-1)^2-2(2t-1)-1 Please Show Steps.

OpenStudy (sasogeek):

ok umm if you have \( (a-b)^2\) it is the same as \( (a^2-2ab+b^2) \) so we simply have to apply this here and then simplify... \[ (2t-1)^2-2(2t-1)-1 \]\[=( \ (2t)^2-2(2t)+1\ \ )-(4t-2)-1\]\[ =4t^2-4t+1-4t+2-1\]\[=4t^2-8t+2\] hence \(f(2t-1)=4t^2-8t+2 \)

OpenStudy (anonymous):

the answer is simple based on (a+b)^2=a^2+b^2+2ab; so the answer would be (2t-11I^2=4t^2+1-4t and then 2(2t-1)=4t-2 so over all goes like this 4t^2+1-4t-(4t-2)-1=4t^2+1-4t-4t+2-1=4t^2-8t+2